OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to y from the second equation is substituted into the first equation. 2x – 7y = 4 3x + y = –17

OpenStudy (anonymous):

2(3x – 17) – 7y = 4 2(–3x – 17) – 7y = 4 2x – 7(3x – 17) = 4 2x – 7(–3x – 17) = 4

OpenStudy (anonymous):

choices above ^^^

OpenStudy (anonymous):

D

OpenStudy (anonymous):

B

OpenStudy (anonymous):

hmmmm